### v1

@@ Abstract
Despite recent strides made by AI in image processing, the issue of mixed exposure, pivotal in many real-world scenarios like surveillance and photography, remains inadequately addressed. Traditional image enhancement techniques and current transformer models are limited with primary focus on either overexposure or underexposure. To bridge this gap, we introduce the Unified-Exposure Guided Transformer (Unified-EGformer). Our proposed solution is built upon advanced transformer architectures, equipped with local pixel-level refinement and global refinement blocks for color correction and image-wide adjustments. We employ a guided attention mechanism to precisely identify exposure-compromised regions, ensuring its adaptability across various real-world conditions. U-EGformer, with a lightweight design featuring a memory footprint (peak memory) of only $\sim$1134 MB (0.1 Million parameters) and an inference time of 95 ms (9.61x faster than the average [^1] Average inference time is calculated among representative models in Table 2. in Ye et al., [67]), is a viable choice for real-time applications such as surveillance and autonomous navigation. Additionally, our model is highly generalizable, requiring minimal fine-tuning to handle multiple tasks and datasets with a single architecture.
Keywords:
Computer Vision Image Processing Image Restoration Low-Light Image Enhancement Unified Learning
@@ 1 Introduction
Figure 1: Sub-figures [a,b,c] show the handcrafted mixed exposure dataset by Zheng et al. [74]; [d-i][4] illustrate real-world scenarios of underexposure, overexposure, and mixed-exposures. [j,k] demonstrates the problem practically.
@@
AI-driven image processing have significantly broadened the scope for enhancing visual media quality. A critical challenge within the low-light image enhancement (LLIE) domain is addressing mixed exposure in images [Fig. 1 g.], where a single frame contains both underlit (below 5 lux, including underexposed)222insufficient brightness in an image where details are lost due to lack of signal. and overlit (overexposed)333excessive brightness in an image where details are lost due to signal clipping or saturation. regions. This issue extends beyond academic interest and has significant real-world implications. For instance, in video calls (e.g., in cafeterias; [Fig. 1 e.]) and live streaming, low-light enhancement is pivotal for clear visual communication. Other areas of application include autonomous driving, surveillance and security, photography, etc. Professional photographers often use high-end DSLR cameras and meticulously adjust settings such as aperture, ISO, and utilize specialized filters to mitigate exposure issues. However, such pre- and post-processing techniques are often not practical.
@@
Existing methods for correcting mixed exposure images have typically treated underexposure ([15, 47, 55]) and overexposure as separate challenges within the low-light image enhancement task. Although these methods (such as FECNet [19], ELCNet+ERL [20], IAT [11]) have made progress, they commonly assume uniform scene illumination, leading to global adjustments that either brighten or darken the entire image. Such approaches falls short when dealing with images that have both overexposed and underexposed regions due to non-uniform lighting, resulting in suboptimal performance. For example, ZeroDCE [15] and RUAS [37] can worsen overexposure in background regions while trying to enhance underexposed foreground subjects [Fig. 1 k.].
@@
We see this as a gap in the literature that hasn’t been fully addressed, despite some efforts such as LCDNet [54] and night enhancement approaches [28, 4, 2]. The motivation for our work arises from the need to address these limitations by developing a solution that can handle mixed exposure scenarios effectively and is suitable for deployment on edge devices. A predominant challenge when applying a low-light enhancement approach to a multi-exposed image is when a model trained solely on underexposed paired images inadvertently exacerbates overexposed regions [Fig. 1 k.], and vice versa. Consequently, the issue of mixed exposure emerges as a pivotal yet largely unexplored frontier. Notably, mobile phone cameras face this issue acutely, requiring lightweight, low-latency models that can operate within the device’s resource constraints while delivering high-quality image enhancement. Our goal is to develop a solution that not only addresses these issues but also offers faster inference and a lower memory footprint, making it ideal for practical applications on edge devices.
@@
This paper aims to develop an effective and computing-efficient approach to tackle mixed exposure in low-light image enhancement. Our major contributions are described as follows:
@@
    We introduce a unified novel framework within the transformer architecture that leverages attention- and illuminance-maps to enhance precision in processing affected regions on a pixel-level, addressing the challenges of underexposure, overexposure, and mixed exposure in images as a single task.
@@
    Our method achieves remarkable efficiency, with an average inference speed of 0.095 seconds per image444computed over LOL-v2 test dataset following previous benchmarks [11, 15, 8, 26]., significantly faster with lesser memory consumption than many existing frameworks. Coupled with a compact architecture of only 101 thousand parameters, the model is ideal for deployment on resource-constrained edge devices.
@@
    We present a novel “Multiplicative-Additive” loss function that intelligently combines contrast scaling and brightness shifting to adaptively enhance images, improving dynamic range and preventing over-smoothing.
@@ 2 Related Work
Traditional to Advanced Deep Learning Techniques. In the realm of image enhancement and exposure corrections, significant strides have been made to address the challenges posed by exposure scenarios. Early techniques [5, 23, 33] leveraged contrast-based histogram equalization (HE), laying the groundwork for more advanced methods. These initial approaches were followed by studies in Retinex theory, which focused on decomposing images into reflection and illumination maps [32, 31]. The advent of deep learning transformed exposure correction, with a shift from enhancing low-light images to addressing both underexposure and overexposure [1, 2, 9, 15, 55, 63, 69, 73]. The notable work of Afifi et al. [1] stands out, employing deep learning to simultaneously address underexposure and overexposure, a task not adequately tackled by previous methodologies. There was a momentous shift towards convolutional neural network (CNN)-based methods, achieving state-of-the-art results and improving the accuracy and efficiency of exposure correction algorithms [15, 29, 47, 51, 55, 64].
@@
Addressing the Challenges of Mixed Exposure. Despite these advancements, the challenges of mixed exposure have remained relatively unaddressed in high-contrast scenarios. Benchmark datasets such as LOL [60], LOL-[4K;8K] [57], SID [7], SICE [4], and ELD [61] offer limited mixed exposure instances, highlighting a gap in both data and models. Synthetic datasets like SICE-Grad and SICE-Mix are enabling the development of methods tailored to mixed exposure scenarios [38, 75, 12, 50]. Some representative methods of deep learning like RetinexNet [60] and KIND [73], focusing on illumination and reflectance component restoration in images, achieved good performance, most methods focus on either underexposure or overexposure correction, failed to correct various exposures. More recent studies have focused on addressing the challenges of correcting both underexposed and overexposed images [54, 66], a task complicated by the differing optimization processes required for each type of exposure. MSEC [2], a revolutionary work in this area, utilizes a Laplacian pyramid structure to incrementally restore brightness and details, dealing with a range of exposure levels.
@@
To manage the correction of a wide range of exposures, several recent works were proposed, such as Huang et al. [17] using exposure normalization, CMEC [48] using exposure-invariant spaces to represent exposures, and ECLNet [17] using bilateral activation for exposure consistency. Moreover, Wang et al. [54] tackle the issue of uneven illumination, while FECNet [18] uses a Fourier-based approach for lightness and structure. Still challenges remain unsolved: (1) Handling nonuniform illumination, (2) Simultaneously addressing both overexposure and underexposure within the same frame, and (3) Ensuring that global adjustments do not adversely affect local regions. Our work addresses these challenges by introducing a unified framework that uses attention and illuminance maps to process mixed exposure regions more precisely.
[…]
Despite these obstacles, the transformer based development for image enhancement has propelled forward progress in the field. Methods like ViT [13] MobileViT [46], CVNet [45], IAT [11], LLformer [57], [21] have shown substantial promise. It also applied to other sub-fields of image enhancements like image dehazing [52, 35], super-resolution [41, 10], image-denoising [65].
@@ 3 Methodology: Unified-EGformer
Unifed-EGformer achieves image enhancement through an Attention Map Generation mechanism that identifies exposure adjustment regions, a Local Enhancement Block for pixel-wise refinement, a Global Enhancement Block for color and contrast adjustment, and an Exposure Aware Fusion (EAF) block that fuses features from both enhancements for balanced exposure correction Fig. 2.
@@ 3.1 Guided Map Generation
Thresholding.  To highlight the sub-regions of the images with impacted exposure problems, we need a way to point out those impacted set of pixels within the input. We use Otsu thresholding, a traditional yet effective technique. It is a global thresholding technique for automatic thresholding that works by selecting the threshold to minimize intra-class variance (variance within a class) or maximize inter-class variance (variance between classes). However, this method induces granular noise in the image [49] due to the non-uniform pixels in low lux regions. The noise is highlighted by the resultant masks as shown in Fig. 3, and will influence the subsequent exposure correction.
To mitigate noise, we implemented adaptive thresholding using pixel blocks and downsampled images. We further reduced noise creep with nearest neighbor downsampling and Gaussian blur. Integrating Charbonnier loss [6] into our attention map mechanism encouraged smoother transitions in areas of high gradient variance, specifically targeting denoising. This component, combined with the SSIM loss that is applied directly on the input, synergistically contributes to noise reduction.
@@
Our implementation of threshold selection is as follows. First, we calculate the Otsu average across the training set to establish a baseline for automatic thresholding. We then apply this average threshold to each image in the dataset. Using a data set-specific threshold, this method ensures a more uniform application555reducing noise propensity in low-light conditions and enhancing exposure correction consistency of the Otsu method.
Attention Map Generator.
The Unified-EGformer begins with a guided attention map generator, designed to identify regions within an image affected by mixed exposure. This process involves generating a map $\boldsymbol{M_{g}}\in\mathbb{R}^{B\times H\times W\times C}$, where $H$, $W$, and $C$ represent the height, width, and channel dimensions of the input image $x\in\mathbb{R}^{H\times W\times C}$. This map, $m\in\boldsymbol{M_{g}}$, is used in an element-wise dot product with the image, resulting in a guided input image that undergoes underexposed, overexposed, or mixed exposure enhancement, as depicted in Fig. 4, demonstrating how we apply Otsu thresholding to get attention masks labels.
The architecture incorporates improved Swin transformer [39] blocks, leveraging Axis-based Multi-head Self-Attention (A-MSA) and Dual Gated Feed-Forward Network (DGFN) [57, 53, 13] for efficient, fast and focused feature processing. The A-MSA reduces computational load by applying self-attention across height and width axes sequentially, optimizing for local contexts within high-resolution images, as illustrated in Fig. 5 (see details in Supplement Section 7.2). The DGFN introduces a dual gating mechanism to the feedforward network, allowing selective emphasis on critical features necessary to distinguish and correct underexposed and overexposed areas effectively (see Supplement Section 7.3).
@@
Figure 6: Ground-truth (a) with it’s corresponding illuminance map [b], inverse illuminance maps [c, and the enhanced results showcasing the effectiveness of the inverse illuminance normalization [d]. Note that the illuminance maps are using false color.
We incorporate the generation of illumination map $I_{illum}$ into the global enhancement block providing a foundational layer for exposure correction. Unlike the complex attention mechanisms required for local block, generating an illuminance map leverages a direct conversion from RGB to luminance ($I_{rgb}\rightarrow I_{illum}$), offering a simpler and faster solution ($I_{illum}=\mathbf{W}\times I_{rgb}$), where $\mathbf{W}$ corresponds to the luminosity method [^6] RGB to Y-channel: https://www.itu.int/dms˙pubrec/itu-r/rec/bt/R-REC-BT.601-7-201103-I!!PDF-E.pdf. The block can utilize the illuminance information to dynamically adjust global parameters such as color balance and exposure levels, ensuring the enhancements are computationally efficient. Our ablation studies Tab .8 confirms that these enhancements are perceptually meaningful compared to the baseline without the illumination map.
@@ 3.2 Unified-Enhancement
Local Enhancement Block (LEB).
The LEB takes a dot product of attention map $\mathcal{A}(x)$ and the sRGB $I(x)$ image, that uniquely forms a objective mapping from input to output. LEB applies a lightweight convolutional block inspired from DETR, PE-Yolo, UWFormer localized pixel-wise refinement. We utilize a adaptive instance based color normalization for capturing a wider range of colors based on individual inputs. Unlike traditional methods, our approach maintains the integrity of features without down-sampling and up-sampling. The output of this block calculates multiplicative ($M_{L}$) and additive ($A_{L}$) correction factors through a feed-forward network, enabling precise pixel-wise enhancement. The local enhancement is formulated as:
@@
Global Enhancement Block (GEB).
Complementary to the LEB, the global enhancement block adjusts the image’s overall exposure through adaptive gamma correction and color balance. Unlike static bias adjustments [11], this block dynamically calculates the gamma correction factor and color transformation parameters based on the image’s content. We implement a convolutional subnetwork with GELU activations and adaptive average pooling, followed by a sigmoid to automatically compute global parameters, denoted by $\theta$. This adaptive approach to global enhancement allows for a more nuanced and content-aware balance of contrast and color. The global correction function is described as:
$G(I)=f_{\gamma}(I;\theta),\quad I\in\mathbb{R}^{H\times W\times C}$ (2)

### v2

@@ Abstract
Despite recent strides made by AI in image processing, the issue of mixed exposure, pivotal in many real-world scenarios like surveillance and photography, remains a challenge. Traditional image enhancement techniques and current transformer models are limited with primary focus on either overexposure or underexposure. To bridge this gap, we introduce the Unified-Exposure Guided Transformer (Unified-EGformer). Our proposed solution is built upon advanced transformer architectures, equipped with local pixel-level refinement and global refinement blocks for color correction and image-wide adjustments. We employ a guided attention mechanism to precisely identify exposure-compromised regions, ensuring its adaptability across various real-world conditions. U-EGformer, with a lightweight design featuring a memory footprint (peak memory) of only $\sim$1134 MB (0.1 Million parameters) and an inference time of 95 ms (over 9x faster than typical existing implementations, is a viable choice for real-time applications such as surveillance and autonomous navigation. Additionally, our model is highly generalizable, requiring minimal fine-tuning to handle multiple tasks and datasets with a single architecture.
@@ 1 Introduction
Figure 1: Sub-figures [a,b,c] show the handcrafted mixed exposure dataset by Zheng et al. [zheng2022low]; images [d-i] from Cai et al. [Cai2018deep] illustrate real-world scenarios of underexposure, overexposure, and mixed-exposures. Images [j,k] demonstrate the problem practically.
@@
AI-driven image processing has significantly broadened the scope for enhancing visual media quality. A critical challenge within low-light image enhancement is addressing mixed exposure in images as in Fig. 1 g., where a single frame contains both underlit111insufficient brightness in an image where details are lost due to lack of signal. (below 5 lux, including underexposed) and overlit222excessive brightness in an image where details are lost due to signal clipping or saturation. (overexposed) regions. This issue extends beyond academic interest and has significant real-world implications. For instance, in video calls (e.g., in cafeterias as in Fig. 1 e.) and live streaming, low-light enhancement is pivotal for clear visual communication. Other areas of application include autonomous driving, surveillance and security, photography, etc. Professional photographers often use high-end DSLR cameras and meticulously adjust settings such as aperture, ISO, and utilize specialized filters to mitigate exposure issues. However, such pre- and post-processing techniques are often not practical.
@@
Existing methods for correcting mixed exposure images have typically treated underexposure [guo2020zero, moran2020deeplpf, Wang_2019_CVPR] and overexposure as separate challenges within the low-light image enhancement task. Although these methods (such as FECNet [10.1145/3343031.3350855], ELCNet+ERL [huang2023learning], IAT [Cui_2022_BMVC]) have made progress, they commonly assume uniform scene illumination, leading to global adjustments that either brighten or darken the entire image. Such approaches falls short when dealing with images that have both overexposed and underexposed regions due to non-uniform lighting, resulting in suboptimal performance. For example, ZeroDCE [guo2020zero] and RUAS [ll_benchmark] can worsen overexposure in background regions while trying to enhance underexposed foreground subjects as shown in Fig. 1 k.
@@
We see this as a gap in the literature that has not been fully addressed, despite some efforts such as LCDNet [wang2022local] and night enhancement approaches [jin2022unsupervised, Cai2018deep, afifi2021learning]. The motivation for our work arises from the need to address these limitations with a solution that can handle mixed exposure scenarios effectively and is suitable for deployment on edge devices. An additional challenge when applying a low-light enhancement approach to a multi-exposed image is when a model trained solely on underexposed paired images inadvertently exacerbates overexposed regions, as seen in Fig. 1 k, and vice versa. Consequently, the issue of mixed exposure emerges as a pivotal yet largely unexplored frontier. Notably, mobile phone cameras face this issue acutely, requiring lightweight, low-latency models that can operate within the device’s resource constraints while delivering high-quality image enhancement. Our goal is to develop a solution that not only addresses these issues but also offers faster inference and a lower memory footprint, making it ideal for practical applications on edge devices.
@@
This paper develops an effective and computing-efficient approach to tackle mixed exposure in low-light image enhancement. Our major contributions are as follows:
@@
    We introduce a novel unified framework within the transformer architecture that leverages attention- and illuminance-maps to enhance precision in processing affected regions at the pixel-level, addressing the challenges of underexposure, overexposure, and mixed exposure in images as a single task.
@@
    Our method achieves remarkable efficiency, with an average inference speed of 0.095 seconds per image333computed over LOL-v2 test dataset following previous benchmarks [Cui_2022_BMVC, guo2020zero, chen2021pre, jiang2017learning]., significantly faster with lesser memory consumption than many existing frameworks. Coupled with a compact architecture of only 101 thousand parameters, the model is ideal for deployment on edge devices.
@@
    We present a novel “MUL-ADD” loss function that intelligently combines contrast scaling and brightness shifting to adaptively enhance images, improving dynamic range and preventing over-smoothing.
@@ 2 Related Work
Traditional to Advanced Deep Learning Techniques. In the realm of image enhancement and exposure correction, significant strides have been made to address the challenges posed by exposure scenarios. Early techniques [celik2011contextual, ibrahim2007brightness, lee2013contrast] leveraged contrast-based histogram equalization (HE), laying the groundwork for more advanced methods. These initial approaches were followed by studies in Retinex theory, which focused on decomposing images into reflection and illumination maps [land1986alternative, land1977retinex]. The advent of deep learning transformed exposure correction, with a shift from enhancing low-light images to addressing both underexposure and overexposure [afifi2020deep, afifi2021learning, chen2018deep, guo2020zero, Wang_2019_CVPR, xu2020learning, yu2018deepexposure, 10.1145/3343031.3350926]. The notable work of Afifi et al. [afifi2020deep] stands out, employing deep learning to simultaneously address underexposure and overexposure, a task not adequately considered by previous methodologies. There was a momentous shift towards convolutional neural network (CNN)-based methods, achieving state-of-the-art results and improving the accuracy and efficiency of exposure correction algorithms [guo2020zero, kim2021representative, moran2020deeplpf, sharma2021nighttime, Wang_2019_CVPR, xu2022snr].
@@
Addressing the Challenges of Mixed Exposure. Despite these advancements, the challenges of mixed exposure have remained relatively unaddressed in high-contrast scenarios. Benchmark datasets such as LOL [WeiWY018], LOL-[4K;8K] [wang2023ultra], SID [ChenCXK18], SICE [Cai2018deep], and ELD [wei2020physics] offer limited mixed exposure instances, highlighting a gap in both data and models. Synthetic datasets like SICE-Grad and SICE-Mix are enabling the development of methods tailored to mixed exposure scenarios [liu2021benchmarking, zheng2020optical, cui2021multitask, sasagawa2020yolo]. Some representative methods of deep learning like RetinexNet [WeiWY018] and KIND [10.1145/3343031.3350926], focusing on illumination and reflectance component restoration in images, achieved good performance, but most methods emphasize either underexposure or overexposure correction, and fail to correct the combination. More recent studies have attempt to address the challenges of correcting both underexposed and overexposed images [wang2022local, drbn], a task complicated by the differing optimization processes required for each type of exposure. MSEC [afifi2021learning], a revolutionary work in this area, utilizes a Laplacian pyramid structure to incrementally restore brightness and details, dealing with a range of exposure levels.
@@
To manage the correction of a wide range of exposures, several recent works were proposed, such as Huang et al. [huang2022exposure] using exposure normalization, CMEC [nsampi2021learning] targeting exposure-invariant feature spaces for exposure consistency, and ECLNet [huang2022exposure] using bilateral activation for exposure consistency. Moreover, Wang et al. [wang2022local] tackle the issue of uneven illumination, while FECNet [huang2022deep] uses a Fourier-based approach for lightness and structure. Still challenges remain unsolved: (1) handling nonuniform illumination, (2) simultaneously addressing both overexposure and underexposure within the same frame, and, (3) ensuring that global adjustments do not adversely affect local regions. Our work addresses these challenges by introducing a unified framework that uses attention and illuminance maps to process mixed exposure regions more precisely.
Emerging Trends with Computational Challenges. Recent studies have begun addressing the dual challenge of correcting under and overexposed images through innovative architectures, including transformers [dosovitskiy2020image], despite their computational intensity as noted in works like Vaswani et al. [NIPS2017_3f5ee243]. The realm of image enhancement has seen remarkable models with human-level enhancement capabilities, such as ExposureDiffusion [wang2023exposurediffusion], Diff-retinex [yi2023diff], wavelet-based diffusion [jiang2023low], PyDiff (Pyramid Diffusion) [zhou2023pyramid], Global structure aware diffusion [hou2024global], LLDiffusion [wang2023lldiffusion], and Maximal diffusion values [kim2019low], demonstrating significant advancements. However, these resource-intensive models, alongside the emerging vision-language models (VLMs) in image restoration, present deployment challenges on edge devices, often requiring tens of seconds to process a single HD or FHD image [Wang_2023_CVPR, luo2023controlling]. In contrast, Unified-EGformer achieves an average inference speed of $\sim$}200 milliseconds on HD images, $9.6\times$ faster on average inference time among representative models in Tab. 2 in Ye et al. [s24041345].
@@ 3 Methodology: Unified-EGformer
Unifed-EGformer achieves image enhancement through an Attention Map Generation mechanism that identifies exposure adjustment regions, a Local Enhancement Block for pixel-wise refinement, a Global Enhancement Block for color and contrast adjustment, and an Exposure Aware Fusion (EAF) block that fuses features from both enhancements for balanced exposure correction as shown in Fig. 2.
@@ 3.1 Guided Map Generation
Thresholding. To highlight the sub-regions of the images with impacted exposure problems, we need a way to point out those impacted set of pixels within the input. We use Otsu thresholding [otsu], a traditional yet effective technique. It is a global thresholding technique for automatic thresholding that works by selecting the threshold to minimize intra-class variance (variance within a class) or maximize inter-class variance (variance between classes). However, this method induces granular noise in the image [otsu] due to the non-uniform pixels in low lux regions. The noise is highlighted by the resultant masks as shown in Fig. 3, and will influence the subsequent exposure correction.
To mitigate noise, we implemented adaptive thresholding using pixel blocks and downsampled images. We further reduced noise creep with nearest neighbor downsampling and Gaussian blur. Integrating Charbonnier loss [charbonnier] into our attention map mechanism encouraged smoother transitions in areas of high gradient variance, specifically targeting denoising. This component, combined with the SSIM loss that is applied directly on the input, synergistically contributes to noise reduction.
@@
Our implementation of threshold selection is as follows. First, we calculate the Otsu average across the training set to establish a baseline for automatic thresholding. We then apply this average threshold to each image in the dataset. Using a data set-specific threshold, this method ensures a more uniform application444reducing noise propensity in low-light conditions and enhancing exposure correction consistency of the Otsu method.
Attention Map Generator. The Unified-EGformer begins with a guided attention map generator, designed to identify regions within an image affected by mixed exposure. This process involves generating a map $\boldsymbol{M_{g}}\in\mathbb{R}^{B\times H\times W\times C}$, where $H$, $W$, and $C$ represent the height, width, and channel dimensions of the input image $x\in\mathbb{R}^{H\times W\times C}$. This map, $m\in\boldsymbol{M_{g}}$, is used in an element-wise dot product with the image, resulting in a guided input image that undergoes underexposed, overexposed, or mixed exposure enhancement, as depicted in Fig. 4, demonstrating how we apply Otsu thresholding to get attention masks labels.
The architecture incorporates improved Swin transformer [Liu_2021_ICCV] blocks, leveraging Axis-based Multi-head Self-Attention (A-MSA) and Dual Gated Feedforward Network (DGFN) [wang2023ultra, NIPS2017_3f5ee243, dosovitskiy2020image] for efficient, fast and focused feature processing. The A-MSA reduces computational load by applying self-attention across height and width axes sequentially, optimizing for local contexts within high-resolution images, as illustrated in Fig. 5 (see details in Supplement Section 7.2The DGFN introduces a dual gating mechanism to the feedforward network, allowing selective emphasis on critical features necessary to distinguish and correct underexposed and overexposed areas effectively (see Supplement Section 7.3).
@@
Figure 6: Ground-truth (a,e) with it’s corresponding illuminance map [b,f], inverse illuminance maps [c,g], and the enhanced results showcasing the effectiveness of the inverse illuminance normalization [d,h]. Note that the illuminance maps are using false color.
We incorporate the generation of an illumination map $I_{\text{illum}}$ into the global enhancement block, providing a foundational layer for exposure correction. Unlike the complex attention mechanisms required for a local block, generating an illumination map leverages a direct conversion from RGB to luminance ($I_{\text{rgb}}\rightarrow I_{\text{illum}}$), offering a simpler and faster solution ($I_{\text{illum}}=\mathbf{W}\times I_{\text{rgb}}$), where $\mathbf{W}$ corresponds to the luminosity method [series2011studio]. The block can utilize the illuminance information to dynamically adjust global parameters such as color balance and exposure levels, ensuring the enhancements are computationally efficient. Our ablation studies shown in Tab. 3 confirm that these enhancements are perceptually meaningful compared to the baseline without the illumination map.
@@ 3.2 Unified-Enhancement
Local Enhancement Block (LEB). The LEB takes a dot product of attention map $\mathcal{A}(x)$ and the sRGB $I(x)$ image, that uniquely forms an objective mapping from input to output. The LEB applies a lightweight convolutional block inspired from PE-Yolo [yin2023peyolopyramidenhancementnetwork], UWFormer [chen2024uwformerunderwaterimageenhancement] localized pixel-wise refinement. We utilize an adaptive instance-based color normalization for capturing a wider range of colors based on individual inputs. Unlike traditional methods, our approach maintains the integrity of features without down-sampling and up-sampling. The output of this block calculates multiplicative ($M_{L}$) and additive ($A_{L}$) correction factors through a feed-forward network, enabling precise pixel-wise enhancement. The local enhancement is formulated as:
@@
Global Enhancement Block (GEB). Complementary to the LEB, the global enhancement block adjusts the image’s overall exposure through adaptive gamma correction and color balance. Unlike static bias adjustments [Cui_2022_BMVC], this block dynamically calculates the gamma correction factor and color transformation parameters based on the image’s content. We implement a convolutional subnetwork with GELU activations and adaptive average pooling, followed by a sigmoid to automatically compute global parameters, denoted by $\theta$. This adaptive approach to global enhancement allows for a more nuanced and content-aware balance of contrast and color. The global correction function is described as:
$\text{G}(I)=f_{\gamma}(I;\theta),\quad I\in\mathbb{R}^{H\times W\times C}$ (2)